OpenStudy (anonymous):

How would I solve this? 5x+4=6x-3

OpenStudy (anonymous):

You substract 5x from both sided os the equasion and add 3 to both of them. Thus you receive 3+4=6x-5x=x thus x=7

OpenStudy (mathstudent55):

5x + 4 = 6x - 3 Subtract 5x from both sides: 5x + 4 = 6x - 3 -5x -5x ----------------- 4 = x - 3 Add 3 to both sides: 4 = x - 3 +3 +3 ------------- 7 = x The solution is x = 7